A voltaic cell is constructed that uses the following reaction and operates at 298 K.

Zn(s) + Ni2+(aq) Zn2+(aq) + Ni(s)

(a) What is the emf of this cell under standard conditions?

___________ V

(b) What is the emf of this cell when [Ni2+] = 2.88 M and [Zn2+] = 0.187 M?

___________ V

c) What is the emf of the cell when [Ni2+] = 0.129 M and [Zn2+] = 0.961 M?

____________ V
Answer #1

Dear student,

Zn(s) + Niat (ap) 2, 2+ (aq) + Ni (s) Half cell sreaction of given reaction oxidation studion acuan at anode Zna) Zn mg + 20Ecell = 0.5041 Gell- 0.31 + Grosar log o ab eo = 0.51 + 0.02 245 = 0.51+ (0.02955 X 0.872) = 0·51 + 0.025 TE call = 0. 5350

Please give a positive ratings to this answer.

With regards...